### v2

# UltraVSR: Achieving Ultra-Realistic Video Super-Resolution with Efficient One-Step Diffusion Space

Abstract.
Diffusion models have shown great potential in generating realistic image detail. However, adapting these models to video super-resolution (VSR) remains challenging due to their inherent stochasticity and lack of temporal modeling. Previous methods have attempted to mitigate this issue by incorporating motion information and temporal layers. However, unreliable motion estimation from low-resolution videos and costly multiple sampling steps with deep temporal layers limit them to short sequences. In this paper, we propose UltraVSR, a novel framework that enables ultra-realistic and temporally-coherent VSR through an efficient one-step diffusion space. A central component of UltraVSR is the Degradation-aware Reconstruction Scheduling (DRS), which estimates a degradation factor from the low-resolution input and transforms the iterative denoising process into a single-step reconstruction from low-resolution to high-resolution videos. To ensure temporal consistency, we propose a lightweight Recurrent Temporal Shift (RTS) module, including an RTS-convolution unit and an RTS-attention unit. By partially shifting feature components along the temporal dimension, it enables effective propagation, fusion, and alignment across frames without explicit temporal layers. The RTS module is integrated into a pretrained text-to-image diffusion model and is further enhanced through Spatio-temporal Joint Distillation (SJD), which improves temporally coherence while preserving realistic details. Additionally, we introduce a Temporally Asynchronous Inference (TAI) strategy to capture long-range temporal dependencies under limited memory constraints. Extensive experiments show that UltraVSR achieves state-of-the-art performance, both qualitatively and quantitatively, in a single sampling step. Code is available at https://github.com/yongliuy/UltraVSR.
Video super-resolution, Diffusion model, Temporal consistency
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755117ccs: Computing methodologies Reconstructionccs: Computing methodologies Image processingccs: Computing methodologies Computational photography
Figure 1. Quantitative and qualitative VSR comparisons. Top-left: An example on a real-world low-resolution (LR) video frame; Top-right: Quantitative comparison with state-of-the-art VSR approaches on the VideoLQ benchmark (Chan et al., 2022b); Bottom: Visual comparison against both SISR approaches and VSR approaches across three consecutive frames. Our one-step (s1) approach achieves leading scores across all three metrics with visually realistic and temporally coherent results.
1. Introduction
Video super-resolution (VSR) aims to reconstruct high-resolution (HR) videos from low-resolution (LR) inputs, a task that has garnered significant attention in multimedia and computer vision communities. Unlike single-image super-resolution (SISR), VSR presents additional challenges due to the necessity of maintaining temporal consistency across frames while recovering high-frequency details. With the rapid advancement of deep learning, numerous VSR approaches (Chan et al., 2021, 2022a; Fuoli et al., 2019; Gupta et al., 2021; Jin et al., 2023; Leng et al., 2022; Xiao et al., 2024; Ai et al., 2024) have emerged over the past decade. However, these methods still struggle to restore realistic textures under complex and unknown real-world degradations.
Figure 2. Comparison of runtime and metrics among diffusion-based VSR approaches. Note that the runtime is measured with the HR sequence upscaled to 720$\times$1280. Our approach achieves the fastest speed with better quality.
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have achieved remarkable success in image synthesis (Rombach et al., 2022; Dhariwal and Nichol, 2021), image editing (Kawar et al., 2023; Yang et al., 2023) and SISR (Wang et al., 2023a, 2024b). However, adapting these models to VSR remains challenging due to the lack of temporal modeling capabilities. Additionally, the inherent randomness of the diffusion process further exacerbates temporal artifacts such as flickering across frames, e.g., the results of SUPIR (Yu et al., 2024) and OSEDiff (Wu et al., 2024) in Fig. 1. Recent approaches, such as MGLD-VSR (Yang et al., 2025) and Upscale-A-Video (Zhou et al., 2024), have attempted to improve temporal consistency by integrating 3D convolutions, temporal attention mechanisms, and motion propagation modules into image diffusion models. Alternatively, text-to-video diffusion models have also been explored for VSR (Xie et al., 2025). Nevertheless, these approaches suffer from high computational costs due to the use of temporal layers and unreliable motion information estimated from LR videos. Moreover, these approaches require a large number of sampling steps for each frame, making inference inefficient and impractical, especially for memory-constrained processing of high-resolution and long-duration video sequences.
This paper proposes UltraVSR, an efficient one-step diffusion model for real-world VSR. In contrast to multi-step VSR approaches, UltraVSR introduces a Degradation-aware Reconstruction Scheduling (DRS) that reformulates the conventional multi-step iterative denoising paradigm as a single-step reconstruction process. DRS first estimates a degradation factor from the input LR video and then maps it with the diffusion noise schedule, thereby achieving single-step reconstruction from LR input to HR video. This scheme eliminates the randomness in the diffusion process and significantly improves computational efficiency. To maintain temporal consistency across frames, we develop a lightweight yet effective Recurrent Temporal Shift (RTS) module, which enables efficient inter-frame feature propagation, fusion, and alignment through two novel units: (1) RTS-convolution unit, which partitions feature maps by channels and shifts groups temporally across adjacent frames for localized feature alignment using 2D convolution; (2) RTS-attention unit, which incorporates temporal shifting operations on $Key$/$Value$ tensors prior to self-attention computation, allowing the model to capture global contextual dependencies.
During training, the RTS module is embedded into a pretrained text-to-image diffusion model and subsequently optimized through Spatio-temporal Joint Distillation (SJD). SJD introduces dual temporal regularizers to separately model real and synthetic video distributions. By minimizing both spatial and temporal gradient discrepancies between two distributions, SJD effectively guides the one-step reconstruction process toward producing visually realistic and temporally consistent HR videos. At inference, we further propose a memory-efficient Temporally Asynchronous Inference (TAI) strategy that decouples spatial and temporal processing by segmenting long video sequences into mini-batches, which are processed independently and then merged prior to RTS propagation. In this way, TAI significantly reduces memory overhead while enabling the model to capture long-range temporal dependencies. Experiments demonstrate that UltraVSR achieves better performance with significantly faster inference, as shown in Fig. 1 and Fig. 2.
In summary, the main contributions of this paper are as follows:
  • We propose UltraVSR, an efficient one-step diffusion framework by introducing a Degradation-aware Reconstruction Scheduling (DRS) for high-quality video reconstruction. To the best of our knowledge, this is the first work to tackle the VSR problem in just one sampling step.
  • We develop a lightweight Recurrent Temporal Shift (RTS) that enables effective feature propagation, fusion, and alignment across frames by partially shifting feature components temporally, without relying on explicit temporal layers.
  • We introduce Spatio-temporal Joint Distillation (SJD) to enhance temporally coherence and content realism, and propose Temporally Asynchronous Inference (TAI) for memory-efficient long-range dependency modeling.
  • Extensive experiments show that UltraVSR outperforms state-of-the-art methods in both visual quality and temporal consistency, while significantly accelerating inference speed.
2. Related Work
2.1. Real-World VSR
Numerous VSR methods (Chan et al., 2021; Fuoli et al., 2019; Gupta et al., 2021; Li et al., 2023; Liang et al., 2022; Xiao et al., 2020; Leng et al., 2022) have been developed to enhance video quality under real-world degradation. For instance, Chan et al. proposed RealBasicVSR (Chan et al., 2022b), which removes degradations in videos by introducing a pre-cleaning module and a stochastic degradation scheme. More recently, Zhang et al. introduced a channel-attention-based recurrent model, RealViformer (Zhang and Yao, 2024), based on findings from investigating channel and spatial attention in a real-world setting. Despite these advancements, many approaches remain sensitive to complex degradations due to limited modeling capacity, leading to poor generalization.
2.2. Diffusion Model-based VSR
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have demonstrated strong performance in generating realistic details (Rombach et al., 2022; Kawar et al., 2023; Li et al., 2025; Yang et al., 2023, 2024; Blattmann et al., 2023). Recent studies have attempted to scale up diffusion models for the VSR task in two ways. On one hand, Zhou et al.(Zhou et al., 2024) and Yang et al.(Yang et al., 2025) leverage image diffusion models and propose Upscale-A-Video and MGLD-VSR, respectively, by incorporating additional temporal layers and motion information. On the other hand, Xie et al.(Xie et al., 2025) build upon a pretrained text-to-video diffusion model and introduce STAR, which employs ControlNet(Zhang et al., 2023) to inject LR video information into the generation process. However, these approaches often rely on multiple sampling steps and complex temporal layers, resulting in high inference latency. To overcome these challenges, we propose an efficient one-step diffusion model for real-world VSR.
3. Proposed Method
3.1. Overview
Our UltraVSR is built upon the pretrained Stable Diffusion model (Rombach et al., 2022), comprising a VAE encoder/decoder, a text encoder, and a diffusion UNet, as illustrated in Fig.3. To construct an efficient one-step reconstruction model $G_{\theta}$ to generate HR video $V_{HR}$ from LR input $V_{LR}$, we introduce three key components: (i) Degradation-aware Reconstruction Scheduling (DRS): Defines a one-step reconstruction from LR videos to HR videos (Sec. 3.2); (ii) Recurrent Temporal Shift (RTS) Module: Incorporates temporal dependencies into the diffusion model by shifting specific feature components across frames, without relying on explicit temporal layers (Sec. 3.3); (iii) Spatio-temporal Joint Distillation (SJD): Guides training with the content realistic and temporal coherent supervision (Sec. 3.4).
Given an input LR video sequence $V_{LR}$, UltraVSR first estimates a degradation factor $d$ and determines a corresponding reconstruction time step $t_{rec}$. Concurrently, a text prompt is simultaneously extracted from $V_{LR}$ via the text encoder, while a latent representation $z_{LR}$ is derived using the VAE encoder. These variables are fed into the diffusion UNet together to produce intermediate estimates $est.$, which are subsequently refined through DRS to obtain the HR latent representation $z_{HR}$. Finally, $z_{HR}$ is passed to the VAE decoder to produce the high-quality HR output $V_{HR}$.
During training, we incorporate LoRA layers (Hu et al., 2022) into both the VAE encoder and diffusion UNet, achieving efficient parameter tuning while preserving generation priors. To ensure temporal consistency throughout the reconstruction process from $V_{LR}$ to $V_{HR}$, we integrate the RTS module across multiple scales of both the diffusion UNet and VAE decoder, components identified as primary sources of temporal flickering artifacts in the original architecture (Yang et al., 2025; Zhou et al., 2024). SJD is further used to enforce that the one-step reconstruction model generates $V_{HR}$ with both spatial realism and temporally coherence. The detailed network architectures are provided in the supplementary materials. We also propose a Temporally Asynchronous Inference (TAI) strategy for memory-efficient long-range dependency modeling in Sec. 3.5.
Figure 3. Training framework of the proposed UltraVSR.
3.2. Degradation-aware Reconstruction Scheduling
Standard diffusion models (DDPMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015) assume a forward diffusion process that gradually adds noise to real data $\boldsymbol{z}_{0}$ following a fixed noise schedule:
(1) $\boldsymbol{z}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}.$
where $t\in\{0,1,2,..,T\}$, $\bar{\alpha}_{t}=\prod_{i=0}^{t}\alpha_{t}$ is a hyperparameter controlling the noise strength. The reverse diffusion process iteratively denoises from $\boldsymbol{z}{T}$ to $\boldsymbol{z}{0}$, but this inherently results in computationally expensive inference.
Inspired by the consistency models (Song et al., 2023) and one-step generation models (Yin et al., 2024; Nguyen and Tran, 2024; Wang et al., 2024a) that map the point $\mathbf{z}_{t}$ on the diffusion trajectory to the starting point $\mathbf{z}_{0}$ directly, we introduce DRS to enable a one-step reconstruction from LR latent $\mathbf{z}_{LR}$ to HR latent $\mathbf{z}_{HR}$.
As shown in Fig. 3 (a), we first estimate a degradation factor $d\in[0,1]$ to quantify the degradation severity of $V_{LR}$:
(2) $d=DEM(V_{LR}),$
where the degradation estimation module ($DEM$) is implemented using the pretrained CLIP model (Wang et al., 2023a), which is a widely used no-reference quality assessment model. Here, $d=0$ corresponds to maximal video degradation, while $d=1$ indicates minimal degradation. This is consistent with the diffusion process described in Eq. 1: a lower $\bar{\alpha}_{t}$ pushes $\boldsymbol{z}_{t}$ closer to Gaussian noise, whereas a higher $\bar{\alpha}_{t}$ retains stronger fidelity to the real data distribution. Therefore, we formulate a direct mapping between the degradation factor $d$ and the noise schedule by defining $d=\sqrt{\bar{\alpha}_{t}}$, mathematically linking perceptual quality to the denoising process. In this way, the HR latent representation $\mathbf{z}_{HR}$ can be derived as:
(3) $\mathbf{z}_{HR}=\frac{\boldsymbol{z}_{LR}-\sqrt{1-d^{2}}est.}{d},$
where $est.$ denotes the diffusion UNet’s output.
Additionally, we determine the reconstruction time step $t_{rec}$ through nearest-neighbor matching against the predefined noise schedule $\{\sqrt{\bar{\alpha}_{t}}\}_{t=1}^{T}$:
[…]
selecting the time step where $\sqrt{\bar{\alpha}_{t_{rec}}}$ most closely approximates the degradation factor $d$. This mechanism enables efficient one-step reconstruction by leveraging the pretrained generative prior.
3.3. Recurrent Temporal Shift Module
The key to VSR lies in effectively aggregating complementary spatio-temporal information from misaligned frames. To this end, we propose a lightweight yet effective Recurrent Temporal Shift (RTS) module that captures spatio-temporal dependencies without relying on computationally expensive temporal layers. As illustrated in Fig.3 (b), the RTS module performs temporal feature propagation by bidirectionally shifting partitioned features across adjacent frames in a parameter-free manner, inspired by(Lin et al., 2019; Cao et al., 2023; Munoz et al., 2021). It consists of two core components: the RTS-Convolution Unit and the RTS-Attention Unit, which collaborate to enable efficient information exchange across neighboring frames. To improve computational efficiency, we apply channel-reducing and channel-increasing convolutions at the beginning and end of each unit, respectively.
3.3.1. RTS-Convolution Unit.
Given an intermediate feature map $F_{i}=Conv(F_{i}^{input})\in\mathbb{R}^{C\times h\times w}$ for the $i$-th frame, RTS divides it into three equal channel-wise segments: $F_{i}^{-1}$, $F_{i}^{0}$, and $F_{i}^{1}$, where each segment $F_{i}^{j}\in\mathbb{R}^{T\times\frac{c}{3}\times h\times w}$. These segments are then shifted along the temporal dimension with offsets of +1, 0, and -1, respectively, where a positive offset denotes a forward shift, a negative offset denotes a backward shift, and zero implies no shift. The aggregated features for the $i$-th frame are then obtained by concatenating the temporally shifted feature segments:
(5) $F_{i}^{Agg}=Concat(F_{i-1}^{-1},F_{i}^{0},F_{i+1}^{1}).$
This strategy ensures that each frame receives contextual information from both past and future frames. To maintain consistent dimensions throughout the video, zero-padding is applied at the sequence boundaries. The aggregated features are processed via convolutional layers with ReLU activation followed by a residual connection to perform local temporal alignment:
(6) $F_{i}^{RTS-C}=Conv(ReLU(Conv(F_{i}^{Agg})))+F_{i}^{input}.$
3.3.2. RTS-Attention Unit.
Building on the observation that attention-based feature alignment inherently captures strong semantic priors in diffusion models (Cao et al., 2023; Alaluf et al., 2024; Geyer et al., 2023; Tang et al., 2023), we further develop an RTS-attention unit to aggregate high-level global contextual dependencies across frames. Given the $Query$, $Key$, and $Value$ embeddings $\{Q_{i},K_{i},V_{i}\}=Conv(F_{i}^{RTS-C})$ for the $i$-th frame, RTS shifts $K_{i}$ and $V_{i}$ forward and backward temporally, allowing the model to query semantically correlated structures and textures from adjacent frames. We adopt Mutual Self-Attention (MutualSA) (Cao et al., 2023) to aggregate semantic information across frames:
(7) $F_{i}^{MSA}=\sum_{j=-1}^{1}\text{MutualSA}(Q_{i},K_{i+j},V_{i+j}),$
where $\text{MutualSA}(Q_{i},K_{i+j},V_{i+j})=\text{Softmax}(\frac{Q_{i},K_{i+j}^{T}}{\sqrt{r}})V_{i+j}$, $r$ denotes the embedding dimension. The final output of the RTS-Attention Unit is obtained by:
(8) $F_{i}^{RTS-A}=Conv(F_{i}^{MSA})+F_{i}^{RTS-C}.$
Unlike conventional spatial self-attention, which operates independently on each frame, the RTS-attention unit aligns semantic information across the temporal dimension, thereby enhancing the temporal consistency of the reconstructed video.
3.4. Spatio-temporal Joint Distillation
To guide the proposed one-step reconstruction in generating visually realistic and temporally consistent HR videos, we introduce a dual-constrained optimization strategy, termed Spatio-temporal Joint Distillation (SJD). Our SJD shares motivation with other distribution matching generative models, such as SDS (Poole et al., 2022) and VSD (Yin et al., 2024; Wang et al., 2023b), but extends these paradigms from the image domain to video, explicitly modeling inter-frame temporal consistency dependencies.
As illustrated in Fig.3 (c), SJD introduces two Temporal Regularized UNets (TRU), a real TRU $\mathcal{T}_{\text{real}}$ and a fake TRU $\mathcal{T}_{\text{fake}}$. They are derived by integrating our proposed RTS module into the Stable Diffusion model (Rombach et al., 2022). During the training of the one-step reconstruction model $G_{\theta}$, $\mathcal{T}_{\text{real}}$ and $\mathcal{T}_{\text{fake}}$ are dynamically updated using ground-truth video $V_{GT}$ and the synthesized output $V_{HR}$, respectively. To capture both spatial realism and temporal consistency, we convert the noise predictions into video distributions using Eq. 1 at the output of the TRUs. Conceptually, $\mathcal{T}_{\text{real}}$ can be interpreted as providing gradient directions that encourage the reconstruction to become more realistic and temporally coherent. In contrast, $\mathcal{T}_{\text{fake}}$ yields gradients that reflect undesirable characteristics such as visual artifacts and temporal flickering. To this end, we formulate the gradient update rule for $G_{\theta}$ by integrating a content-realistic gradient loss and a temporal-consistency gradient loss:
(9) $\nabla\mathcal{L}_{\text{SJD}}=\nabla\mathcal{L}_{\text{realistic}}+\lambda\nabla\mathcal{L}_{\text{consistency}},$
where $\lambda$ balances content realism and temporal coherence of $V_{HR}$. This encourages the synthesized $V_{HR}$ toward higher perceptual realism and temporally coherence while suppressing visual artifacts and flickering. The content-realistic gradient loss is formulated as:
(10) $\nabla\mathcal{L}_{\text{realistic}}=\mathbb{E}\left[\omega(t)\left(\mathcal{T}_{\text{real}}\left(z_{t},t\right)-\mathcal{T}_{\text{fake}}\left(z_{t},t\right)\right)\frac{dG}{d\theta}\right],$
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, $z_{0}=\mathcal{E}(V_{HR})$, $V_{HR}=G_{\theta}(V_{LR})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, $\mathcal{E}$ denotes the pretrained VAE encoder, $\omega(t)$ is a weighting coefficient to normalize the gradient magnitudes across varying noise levels (Yin et al., 2024). The temporal-consistency gradient loss enhances temporally coherence by penalizing discrepancies in the inter-frame gradients of two TRUs:
(11) $\nabla\mathcal{L}_{\text{consistency}}=\mathbb{E}\Bigg{[}\omega(t)\Big{(}(\mathcal{T}_{\text{real}}(z_{t},t)[:-1]-\mathcal{T}_{\text{real}}(z_{t},t)[1:])$
$-(\mathcal{T}_{\text{fake}}(z_{t},t)[:-1]-\mathcal{T}_{\text{fake}}(z_{t},t)[1:])\Big{)}\frac{dG}{d\theta}\Bigg{]}.$
3.5. Temporally Asynchronous Inference
Existing diffusion-based VSR methods exhibit critical dependencies on temporal modules, incurring prohibitive memory overheads. As a result, such models require video segmentation into short clips for piecewise processing, as shown in Fig. 4 (a). For instance, under 48GB GPU memory constraints (NVIDIA RTX 8000), MGLD-VSR (Yang et al., 2025) processes $\leq$6 frames while STAR (Xie et al., 2025) handles $\leq$15 frames for 2K (1440$\times$2560) reconstruction. This fragmentation disrupts temporal continuity and fundamentally limits long-range dependency modeling. UltraVSR circumvents this limitation through implicit temporal modeling via RTS propagation, eliminating explicit temporal layers. Building on this architectural innovation, we propose a Temporally Asynchronous Inference (TAI) strategy that enables efficient capture of long-range temporal dependencies.
As illustrated in Fig. 4 (b), given an $N$-frames low-resolution video $V_{LR}$, TAI first divides it into $m$ non-overlapping sub-sequences $\{V_{LR_{1}},V_{LR_{2}},\dots,V_{LR_{m}}\}$, each containing $b$ frames. Then, these sub-sequences are processed sequentially through 2D spatial layers to obtain feature maps $\{F_{b_{1}},F_{b_{2}},...F_{b_{m}}\}$, drastically reducing peak memory consumption. At the beginning of the RTS propagation, all these features are concatenated into a unified feature tensor $F_{N}$, which is then augmented by bidirectional RTS propagation to achieve efficient temporal information exchange between frames. The temporally shifted features are then divided back into sub-sequences for further refinement.
As a result, TAI enables memory-efficient inference while maintaining long-term temporal dependencies, capable of processing up to 120 frames at 1440$\times$2560 resolution identical hardware constraints, an 8$\times$ throughput improvement over STAR (Xie et al., 2025). This breakthrough makes UltraVSR particularly suitable for high-quality reconstruction of long-duration video sequences under resource-limited environments.
Figure 4. Previous inference scheme vs. the proposed TAI.
Table 1. Quantitative comparsion with state-of-the-art VSR approaches on both synthetic and real-world benchmarks. The best and second best results are highlighted and underlined, respectively.
Datasets Metrics BasicVSR++ (Chan et al., 2022a) OVSR (Yi et al., 2021) RealBasicVSR (Chan et al., 2022b) VRT (Liang et al., 2024) RealViformer (Zhang and Yao, 2024) Upscale-A-Video (Zhou et al., 2024) MGLD-VSR (Shi et al., 2024) STAR (Xie et al., 2025) Ours REDS4 PSNR 33.77 23.55 28.59 32.97 28.18 25.62 27.49 24.08 24.50 SSIM 0.9173 0.6743 0.8066 0.9007 0.7896 0.6761 0.7783 0.6721 0.6962 MUSIQ 67.00 54.44 67.03 65.68 64.58 32.30 58.94 60.94 70.07 CLIP-IQA 0.3149 0.3383 0.3724 0.3743 0.3558 0.1843 0.2908 0.3047 0.4397 MANIQA 0.6752 0.5887 0.6671 0.6699 0.6545 0.3138 0.5837 0.6123 0.6821 DOVER 0.7182 0.6620 0.7209 0.7067 0.7152 0.4053 0.6717 0.7227 0.7304 SPMCS PSNR 23.39 29.87 22.87 23.43 23.45 24.10 23.43 22.95 23.08 SSIM 0.6516 0.8702 0.6087 0.6543 0.6153 0.6386 0.6372 0.6001 0.6202 MUSIQ 62.57 57.10 69.44 63.44 66.10 56.31 64.08 55.14 69.92 CLIP-IQA 0.4341 0.4597 0.4838 0.4388 0.4432 0.3353 0.4685 0.3888 0.5377 MANIQA 0.6263 0.6057 0.6282 0.6134 0.6148 0.4798 0.5562 0.5640 0.6386 DOVER 0.6711 0.6031 0.7506 0.6404 0.7206 0.5675 0.6431 0.6404 0.7616 UDM10 PSNR 34.68 27.12 31.31 36.03 31.84 29.48 29.50 27.93 27.96 SSIM 0.9406 0.8318 0.8936 0.9501 0.9074 0.8458 0.8744 0.8284 0.8346 MUSIQ 63.26 57.94 66.01 61.38 61.86 39.48 64.54 57.34 66.46 CLIP-IQA 0.5091 0.4338 0.4903 0.4697 0.4356 0.2356 0.5083 0.3843 0.4818 MANIQA 0.5729 0.5593 0.6083 0.5501 0.5699 0.3329 0.5495 0.5219 0.6167 DOVER 0.7947 0.7789 0.8248 0.7709 0.8185 0.5500 0.7851 0.8061 0.8361 YouHQ40 PSNR 23.99 22.67 25.26 25.01 25.73 25.20 24.97 23.76 24.75 SSIM 0.5999 0.5670 0.6857 0.6510 0.6974 0.6865 0.6984 0.6593 0.6954 MUSIQ 49.43 45.54 62.12 41.45 62.27 35.86 56.42 41.47 62.71 CLIP-IQA 0.3798 0.3442 0.4661 0.3879 0.4520 0.2460 0.3960 0.3435 0.4928 MANIQA 0.5145 0.4710 0.5539 0.4622 0.5500 0.3401 0.4717 0.4198 0.5847 DOVER 0.6914 0.6877 0.8855 0.6347 0.8657 0.6216 0.8116 0.8084 0.8872 VideoLQ MUSIQ 37.32 29.89 62.48 37.94 58.66 30.88 55.99 54.59 64.42 CLIP-IQA 0.2766 0.2875 0.3663 0.2763 0.3211 0.1817 0.3385 0.3126 0.4599 MANIQA 0.4099 0.3447 0.5721 0.4051 0.5477 0.3025 0.5004 0.5507 0.5907 DOVER 0.5128 0.5141 0.7645 0.5136 0.7399 0.4493 0.7254 0.7034 0.7785
Figure 5. Visual comparisons on synthetic low-quality videos from SPMCS (Yi et al., 2019) and YouHQ40 (Zhou et al., 2024) datasets.
4. Experiments
4.1. Experimental Settings
4.1.1. Datasets.
Similar to the most VSR approaches (Chan et al., 2022b; Yang et al., 2025; Zhang and Yao, 2024), we combine the training and validation sets from the REDS dataset (Nah et al., 2019) for model training, reserving four sequences (REDS4) for testing. We evaluate UltraVSR on four synthetic datasets (i.e., REDS4, SPMCS (Yi et al., 2019), UDM10 (Tao et al., 2017), and YouHQ40 (Zhou et al., 2024)) and a real-world dataset (i.e., VideoLQ (Chan et al., 2022b)) that contains 50 real-world video sequences with diverse degradations. For both synthetic training and testing datasets, LR-HR video pairs are generated following the degradation pipeline of RealBasicVSR (Chan et al., 2022b).
4.1.2. Training Details.
We implement UltraVSR using the PyTorch framework and the Diffusers (Patrick von Platen and Wolf, 2022) library. The input sequence length is set to 6, with the spatial size cropped to 512 $\times$ 512 resolution. We freeze the original spatial modules of the pretrained Stable Diffusion model while only training the newly introduced LoRA and RTS modules. Adam (Kingma and Ba, 2014) optimizer is employed as the optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We use a combination of MSE loss, perceptual loss, temporal GAN loss (Zhou et al., 2023), and SJD regularization loss to train UltraVSR. The real and fake TRUs are trained by minimizing a standard denoising objective (Ho et al., 2020; Vincent, 2011). The weighting coefficient $\lambda$ is set to 0.5. The learning rate is set to $5\times 10^{-5}$ for all models.
Figure 6. Visual comparisons on real-world low-quality videos from VideoLQ (Chan et al., 2022b)dataset.
4.1.3. Evaluation Metrics.
We evaluate super-resolved video quality using three image-based quality metrics: MUSIQ (Ke et al., 2021), CLIP-IQA (Wang et al., 2023a), and MANIQA (Yang et al., 2022), along with one video-specific quality metric, DOVER (Wu et al., 2023). DOVER assesses both content quality and temporal consistency from both aesthetic and technical perspectives. In addition, the full-reference metrics (i.e., PSNR and SSIM) are also given as a reference. However, it should be noted that they poorly respond to realistic visuals (Blau and Michaeli, 2018; Gu et al., 2022; Ledig et al., 2017; Yu et al., 2024; Jinjin et al., 2020) and can not reliably evaluate the performance of diffusion-based VSR models.
4.2. Evaluation with State-of-the-art Methods
To validate the efficacy of UltraVSR, we conduct both quantitative and qualitative comparisons with state-of-the-art VSR methods: BasicVSR++ (Chan et al., 2022a), OVSR (Yi et al., 2021), RealBasicVSR (Chan et al., 2022b), VRT (Liang et al., 2024), RealViformer (Zhang and Yao, 2024), Upscale-A-Video (Zhou et al., 2024), MGLD-VSR (Shi et al., 2024), and STAR (Xie et al., 2025). Due to the extremely high memory requirements of STAR, we had to limit its output resolution to a maximum of 1440 pixels. In addition, we include comparisons with state-of-the-art SISR approaches (i.e., RealESRGAN (Wang et al., 2021), StableSR (Wang et al., 2024c), SUPIR (Yu et al., 2024), ResShift (Yue et al., 2024) and OSEDiff (Wu et al., 2024)) in the supplementary materials.
4.2.1. Quantitative Comparison.
As demonstrated in Tab. 1, we present the quantitative results with existing state-of-the-art VSR approaches. It can be seen that our UltraVSR consistently achieves the highest score across nearly all non-reference metrics on all synthetic benchmarks, demonstrating its superior performance. Although approaches like BasicVSR++ (Chan et al., 2022a) and VRT (Liang et al., 2024) have better performance in terms of PSNR or SSIM, they usually produce unsatisfactory results on real-world degraded videos (see Sec. 4.2.2), as reflected in their lower no-reference metric scores. As for the real-world VSR dataset VideoLQ (Chan et al., 2022b), UltraVSR outperforms other methods across all non-reference metrics (MUSIQ: +1.94, CLIP-IQA: +0.0936, MANIQA: +0.0186, DOVER: +0.014). These notable improvements highlight UltraVSR’s strong ability to enhance complex and diverse real-world videos while preserving realistic details.
Table 2. Efficiency evaluations among diffusion-based VSR approaches. Both trainable/total parameters are measured for each model. The best results are highlighted.
Methods Sampling step Runtime (s) # Params 720$\times$1280 1440$\times$2560 (M) Upscale-A-Video (Zhou et al., 2024) 30 16.1 32.0 $\sim$ 96.6/746 MGLD-VSR (Shi et al., 2024) 50 31.8 204 156.0/1465 STAR (Xie et al., 2025) 15 14.5 110.2 629.9/2139 Ours 1 0.89 2.67 10.5/1912
4.2.2. Qualitative Comparison.
To further assess the effectiveness of the proposed UltraVSR, we present qualitative comparisons on both synthetic datasets (SPMCS (Yi et al., 2019) and YouHQ40 (Zhou et al., 2024) in Fig. 5) and real-world dataset (VideoLQ (Chan et al., 2022b) in Fig. 6). On synthetic videos, it can be observed that the competing approaches struggle to reconstruct accurate texture (e.g., the results of Upscale-A-Video (Zhou et al., 2024) and STAR (Xie et al., 2025) in the first row) and blurriness (e.g., the results of RealViformer (Zhang and Yao, 2024) and the result of Upscale-A-Video (Zhou et al., 2024) in the second row). In addition, MGLD-VSR (Shi et al., 2024) introduces color artifacts in the reconstructed frames. In contrast, our UltraVSR excels in restoring lost textures and structures under complex degradation scenarios, such as the outlines of statues and animal fur. On real-world videos, UltraVSR demonstrates superior texture reconstruction capabilities compared to competing approaches. For instance, in sequence 004, it not only recovers clear and realistic facial expressions but also faithfully reconstructs the structural details of the surrounding brick wall. In sequence 039, UltraVSR successfully restores intricate architectural details and decorations, demonstrating its robustness and generalization ability under challenging degradation scenarios. More comparisons can be found in the supplementary materials.
Table 3. Ablation study of different configurations. Both trainable/total parameters of the one-step reconstruction model are measured for each model variant. The best and second best results are highlighted and underlined, respectively.
[…]
Figure 7. Temporal consistency analysis across SISR, VSR approaches, and different model variants.
4.2.3. Temporal Consistency.
From a quantitative perspective, we compare the DOVER metric, an effective video quality assessment approach, against competing methods in Tab. 1. Due to the inherent randomness in the diffusion model, the DOVER scores of existing diffusion-based VSR methods are generally low. Benefiting from the proposed one-step diffusion space, our UltraVSR achieves the best score on all the datasets (e.g., +0.011 on the SPMCS dataset and +0.014 on the VideoLQ dataset). In addition, Fig. 7 presents visual comparisons against both SISR and VSR approaches. It can be found that due to the inherent randomness and insufficient temporal modeling, results from SUPIR (Yu et al., 2024) and OSEDiff (Wu et al., 2024) exhibit severe temporal inconsistencies. While existing VSR approaches alleviate this issue, they often sacrifice texture details. In contrast, our ultraVSR successfully reconstructs videos with both fine texture structure and temporal coherence.
4.2.4. Efficiency Evaluations.
To evaluate the efficiency of UltraVSR, we compare its sampling step, runtime, and parameters against other diffusion-based VSR approaches in Tab. 2. we assess computational efficiency at two common resolutions: 720p (720$\times$1280) and 2K (1440$\times$2560). All models were evaluated on a single NVIDIA RTX 8000 GPU. Existing diffusion-based VSR approaches typically require dozens of iterative sampling steps, resulting in prohibitively long inference times, especially for 2K video, making them impractical for real-world deployment. Thanks to the efficient one-step diffusion space, our UltraVSR significantly reduces inference time, taking only 0.06$\times$ and 0.08$\times$ that of the second fastest approach at two resolutions. Moreover, UltraVSR has substantially fewer trainable parameters (just 10.5M) compared to other diffusion-based methods, underscoring its efficiency during the training process.
4.3. Ablation Study
We conduct comprehensive experiments to evaluate the contributions of the key components in UltraVSR. The results are presented in Tab. 3 and Fig. 7. We begin with Model A, which serves as our baseline, constructed by: (1) fixing the diffusion time step to 999, following the one-step generation paradigm established in prior works (Yin et al., 2024; Nguyen and Tran, 2024; Frans et al., 2024; Xie et al., 2024); (2) enabling only the trainable LoRA layers; and (3) replacing our proposed Spatio-temporal Joint Distillation (SJD) with the conventional Variational Score Distillation (VSD) (Yin et al., 2024; Wang et al., 2023b). Due to its inability to dynamically perceive degradation levels, Model A exhibits limited reconstruction quality compared to Model B (with DRS). We incrementally enhance Model B with trainable RTS-convolution and RTS-attention units to form Models C and D, respectively, leading to steadily improved inter-frame consistency. We further evaluate the effectiveness of SJD and TAI by comparing Model D, Model E (with a fixed local sequence of length 10), and our model. It can be seen that SJD significantly enhances temporal coherence by introducing inter-frame difference supervision, leading to improved motion continuity across frames compared to VSD (Yin et al., 2024; Wang et al., 2023b). TAI further boosts performance by enabling long-range temporal modeling.
5. Conclusion
We present UltraVSR, the first real-world VSR framework based on an efficient one-step diffusion space. It introduces Degradation-aware Reconstruction Scheduling (DRS) to transform multi-step denoising with one-step LR-to-HR reconstruction. To maintain temporal consistency, we develop the Recurrent Temporal Shift (RTS) module and Spatio-temporal Joint Distillation (SJD) for realistic and coherent outputs, along with Temporally Asynchronous Inference (TAI) for efficient long-range temporal modeling. Extensive experiments show that UltraVSR delivers favorable performance with significantly accelerated inference speed.
6. Acknowledgments
This work was supported in part by the National Major Science and Technology Projects of China under Grant 2009XJTU0016, in part by the National Natural Science Foundation of China under Grant U22B2049.